In a right angled triangle the hypotenuse is four times as long as the perpendicular drawn to it from the opposite vertex. One of the acute angle is a)15° b)30° c)45° d) none of these

Grade:11

3 Answers

uortoiei
7 years ago
c is the answer euto erouh kjghuf oeifdhiv oasidjfh pofhr vjdfbf  uerhg iknf lfiisrn liedvnnv lrinjf fdsjkfj vdfkjns sdljnf lsdfof ilfeki
Vaibhavi
7 years ago
Sorry but I didn`t get the language plz explain it in english And explain it step by step............... ...............
Sakshi
7 years ago
So let the triangle of sides ab, bc and ac, where ac is the hypotenuse of side 4x and the opposite perpendicular side ab is x . Let the angle you needs to find is y. Sin(y) = opposite / hypotenuse = x / 4x = 1/4 ............ y = sin^-1 ( 1/4 ) which is approximately equal to 75 and 15 . Since, it`s an acuteness angle the answer is 15